The final exam in statistics class were normally distributed with a mean of 63 and standard deviation of 5. Find the 90th percentile

1
Expert's answer
2022-05-22T23:08:59-0400

The formula below is used to compute percentiles of a normal distribution.


"x=\mu+z\sigma"

Using "Z=1.2816" the 90th percentile is


"x=63+1.2816\cdot 5=69.408"